Tutorials:

Classical Mechanics Series: Hooke's Law Tutorial Creator of Tutorial: Robert136

Tutorials:

This is a simple tutoring on Hooke's law. Hooke's law identifies the constant of a given string... or k. k=F(applied)/Delta x Where k denotes the force required to stretch or compress a spring per unit of length.

Tutorials:

Usually K is measured in N/m. Let's keep that as our rule of thumb in solving the following question. The length of a stretched spring is 2m with an applied force of 16N. To identify the k constant for the spring, K=16N/2m=8N/m Therefore the k constant for this spring is 8N/m

Tutorials:

p.s Note that there is also a reference point in spring as well. You would measure the stretch or compression from the point of zero.

Tutorials:

Alright, the next equation concerning Hooke's law is this, E(energy stored in a compressed or stretched spring)=1/2(k)Delta(x) Where, E=potential stored k=Hooke's constant x=stretch, or compression of the spring. With this formula alone you can calculate the energy stored in a spring, or potential energy from a spring being stretched or compressed. Therefore, you look at the problem Calculate the spring constant of a spring if a 15 kg mass is suspended from the spring and it stabilizes with a stretch of 12 cm. First you need to look at the resulting force acting downwards to the gravity source, which is accounted for by (9.8m/s^2)(15kg)=147N Next step is you must convert the 12cm to meters 12cm/100cm/m=0.12m Then you would calculate the k constant L=147N/0.12m=1225N/m This tell us how much force is required to stretch or compress the spring by a meter. Now using the formula E=1/2kx^2 E(potential energy stored in the spring)=1/2(1225N/m)(0.12cm)^2=8.82J After accounting for sig figs of 2, The energy stored in the spring is 8.8J.